OpenStudy (anonymous):

Assuming x > 0, what is the simplified fraction form for 5xsqrt3/sqrt48x^2

OpenStudy (anonymous):

√ 48 x² = √(3 * 16) x² = 4√3 x² -> 5√3 x / 4 √3 x² = 5/ 4x